From statistical studies, we know that for 1981 the supply curve for wheat was approximately as follows: Supply: QS = 1800 + 240P Where price is measured in dollars per bushel and quantities are in millions of bushels per year. These studies also indicate that in 1981 the demand curve for wheat was Demand: QD = 3550 – 266P a. Find the market clearing price and equilibrium quantity of wheat for the year1981. b. What do we mean by surplus and deficit? Show us your answer graphically using the above demand and supply functions c. Show the effects of changes in demand and supply in equilibrium price and equilibrium quantity

1
Expert's answer
2022-06-10T18:32:58-0400

"Q_s=1800+240P,"

"Q_d=3550-266P,"

"Q_s=Q_d,"

"1800+240P=3550-266P,"

"506P=1750,"

"P=3.46,"

"Q_s=Q_d=2630."